# supplementary

### v3

@@ 2.2 AI representatives to recover from low voters turnout
Figure 4a illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts. Strikingly, up to 50% of AI representatives are sufficient for a maximum consistency recovery of 53.3% under voting turnouts lower than 80% when equal shares is used. This significant recovery is despite the very low consistency of human-AI individual choices in complex ballot formats with several options (Figure 2a). The recovery is also confirmed for equal shares with the top-8 winners, which is equal to the number of winners with utilitarian greedy (on average 18.1% lower consistency recovery). However, AI representatives fail to recover consistency with utilitarian greedy as the voting aggregation method.
Figure 4: By representing up to half of the human voters who abstain from voting with AI, a significant lost consistency in the collective choice is recovered in participatory budgeting under fair voting aggregation methods such as equal shares and under low number of alternatives. However, further consistency loss is observed under a large number of alternatives with the standard utilitarian greedy method as the number of AI representatives increases. In contrast, single-choice majoritarian elections with highly consistent human-AI individual choice maximize consistency recovery for the whole spectrum of voters turnout by maximizing the AI representatives for the voters who abstain. Nevertheless, for more than 50% of voters turnout, consistency is fully recovered by representing at least 25% of abstained voters with AI. The consistency loss in the voting outcome by low voters turnout (x-axis) is emulated by removing different ratios of human voters. A recovery of consistency (y-axis) is hypothesized by different ratios of abstained voters represented by AI: 25%, 50%, 75% and 100%. The consistency recovery values are the means of all three large language models (see Figure LABEL:fig:LLM_factorial for the results of each large language model). The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea are shown with two voting aggregation methods: utilitarian greedy and equal shares, with all winners and the top-8 (as many as greedy) for a fairer comparison. (C) US elections. (D) Chord visualizations summarizing how the interplay between voters turnout and AI representation yields voting outcomes with consistency gain or loss. (E) Visualizations summarizing the mechanism of consistency recovery through AI representatives for the voters who abstain.
Figure 4b shows that the recovery of consistency is more robust for both voting aggregation methods under low number of alternatives. Consistency is recovered with up to 75% of abstained voters represented by AI under a maximum of 60% voters turnout, after which 50% of AI representatives are sufficient for consistency recovery.
Under the voting scenario of the US elections that comes with highly consistent human-AI individual choices, consistency is fully recovered for $\geq$25of abstained voters represented by AI under a minimum of 50% voters turnout. However, below 50% of voters turnout, election results are very sensitive and consistency recovery is guaranteed only for 100% of abstained voters represented by AI. More detailed results on mixed populations of humans and AI representatives, including the consistency loss when human voters abstain are shown in Figure LABEL:fig:mixed_populations-abstaining.
2.3 Biases explaining AI inconsistencies in choice and preference transitivity
To explain the origin of (in)consistencies in human-AI choice and different voting input methods, a prediction model is designed for each voting scenario. The personal human traits, mapped to different cognitive biases, become the independent variables and the consistency the dependent variable. The prediction model is built using a neural network (see Section 4.4). The relative importance of the personal human traits that explain the AI consistency is calculated using shapley additive explanations (see Section 4.5).
In the case of participatory budgeting in Figure 5a and 5b, the preference for environmental projects (potential conformity bias [40, 41, 42]) shows an average importance of 26.5% (p=0.0258) in explaining human-AI choice consistency (25% and 12% higher consistency among environmental supporters, with p=0.038 for the actual voting scenario). Political belief and ideology are strongly correlated with the actual two alternatives in the US elections (right/conservative for republicans and left/liberal for democrats), therefore, they strongly explain the consistency of human-AI choice (mean relative importance of 98.7%, p<0.02 and 4.8%, p<0.03 respectively). However, this is not the case in participatory budgeting as it is political belief that rather explains inconsistency (relative importance of -51.2% with p<0.02). This demonstrates that the underlying participatory budgeting choices remain politically neutral to a high extent, despite the fact that the City Idea campaign mainly attracted left-wing voters (42.5% versus 31.5% of right wing voters). In contrast, race and discussion of politics explain inconsistencies in the US elections with a mean relative importance of -48.1% (p<0.04) and -64.7% (p<0.04) respectively. Inconsistency is 33% higher than consistency among white people (p<0.019), while for people who discuss politics inconsistency is 63.2% higher that the ones who do not (p=0.046).
@@ 2.3 Biases explaining AI inconsistencies in choice and preference transitivity
Figure 5: Conformity bias, such as preference for environmental projects, positively contribute to the consistency of human-AI choice. On the other hand, unconscious biases such as race and discussion of politics cause inconsistencies. For US elections though, political belief and ideology strongly contribute to consistency. Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families positively contribute to the consistency of large language models among the voting input methods. The relative importance (x-axis) of the personal human traits (y-axis) to predict the consistency of choice and consistency between voting input methods (single choice vs. cumulative/score) is shown for different large language models and the machine learning model. This is calculated using Shapley Additive Explanations. The (A) actual and (B) mock voting in the participatory budgeting campaign of City Idea is shown along with the (C) US elections. The ‘*’ denotes that the feature is statistically significant (p<0.05) for each of the AI models of Llama2, GPT 3, GPT 3.5 and ML respectively. The chord visualizations summarize the biases that explain the (in)consistencies of (D) human-AI choice and (E) voting input method.
The consistency among the voting input methods is explained by preference towards time discounting factors [41] in participatory budgeting, such as sport and welfare projects. The average relative importance is 25.6% and 17.5% respectively (all with p<0.03), with 48.53%, and 37.5% of higher consistency among supporters in the actual voting (p<0.044). Affect heuristics such as preference for projects that benefit families with children also strongly contributes to the consistency with 58.9% higher consistency among supporters in the actual voting (p<0.032). Figures, LABEL:fig:merged3aLABEL:fig:explaination-top-choice, and Table LABEL:table:error illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.
@@ 3 Discussion
Generative AI voting: an inevitable race to safeguard democracy. Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world. On the one hand, generative AI and large language models are expected to become more open, pervasive and accessible to citizens [43, 15]. AI personal assistants are already part of everyday life [44, 45, 46], with their generative version expected to follow. On the other hand, the mandate of more direct, secure and active participation in decision making for public matters is expected to further scale up electronic voting solutions and digital platforms. For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [47]. As Toomas H. Ilves, former president of Estonia, emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen". In the light of these converging advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations. Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare safeguards for digital democracy.
@@
Fair voting design as a safeguard for generative AI voting. Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting. However, a striking finding of this study, built upon our demonstrated real-world evidence in Aarau, is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences. This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies. The momentum for this win-win paradigm shift is tremendous, which would not be possible otherwise to demonstrate without the real-world empirical evidence based on which this study unfolds: fair voting rules such as the method of equal shares promotes under-represented projects and yield more winning projects with higher proportional representation of citizens’ preferences [39, 22]. Nevertheless, how to scale up these democratic blueprints remains an open question, but AI can play a key role here.
@@
Why fair collective choice is resilient to AI biases and inconsistencies. We provide an explanation of this significant and striking finding. There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [22]. For instance, even with 25-50% of the voters, the elected projects do not significantly change [22]. This is despite the fact that projects may end with very different number of total votes, as low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters. Such projects are expected to be an origin of consistency. Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abstaining. As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind and comes with compelling realism and merit for the validity of our significant findings. This comes in stark contrast to other earlier studies [22, 39] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
Mitigating low voters turnout: AI representation is not enough, voting design matters. This perspective is assessed for the first time in this study by emulating voting scenarios with varying voters turnout, while letting voters who abstain to have a representation by generative AI to restore the lost consistency of abstaining. Nevertheless, the inherent inconsistencies of generative AI can hinder the effectiveness of AI representatives. Therefore understanding these opposing dynamics comes with significant merit to mitigate the impact of AI biases and inconsistencies in collective decision making. This is by showing striking evidence that voting design has a determining role on how much consistency can be recovered and what effect this can have for the legitimacy of the voting process and outcomes. Recovering the lost consistency under very low voting turnouts in single-choice majoritarian elections requires populations with very high majorities of AI representatives, which turns out to be illegitimate. In contrast, the lost consistency under higher voting turnouts can be restored even with low proportions of AI representatives. The fair method of equal shares is also here a prominent accelerator for the consistency recovery: fewer AI representatives are sufficient to restore consistency even under very low voter turnouts. It is also evident that such consistency recovery could not be achieved with the standard voting aggregation method.
@@
Trustworthy generative AI voting: a call for research and policy action. What information large language models use to reason about voting decisions is influential for different types biases to manifest. Unconscious and conformity biases, related to political belief, race and environment reinforce human choices or yield inconsistencies. Time discounting biases, related to the support of projects on welfare and sport, explain preference transitivity among different ballot formats. These significant findings demonstrate that training data in generative AI voting are expected to be highly impactful for the representation of the voting outcomes in the population. Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes. For instance, who shall determine the input training data of AI representatives? Should the training data involve only self-determined personal information of voters, or shall these be augmented with more universal knowledge and experts’ opinions? How to protect the privacy and autonomy of voters when training such AI representatives [46]? Will citizens retain power to control AI representatives that reflect their values and beliefs, while remaining accountable? These are some key questions as a basis of a call for action on research and policy in an emerging era of generative AI voting.
@@ 4.1 Emulating AI representatives
We introduce a generalized design to create AI representatives using three large language models and a predictive AI method as a benchmark (referred to as ML) to generate choices (ballots) and emulate AI voting representation.
Input to ballot generation. This comprises of (i) defining the voting scenario, the voting alternatives and the personal human traits of the voter as the input context (LLMs) or features (ML) to inform the ballot choice, (ii) generating the ballots and (iii) aggregating the ballots into a collective decision (see Figure 1).
@@
    Personal human traits: Voter’s information is used to define an AI persona of a human voter, i.e. the AI representative (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2 in SI). This includes (i) socio-demographic characteristics (e.g., gender, age, education, household size), (ii) political interests (e.g., ideological profile, political belief), (iii) personal attitude towards project preferences (e.g., prioritization for green initiatives, sustainable transport, elderly care facilities) and (iv) preferences for the qualities of voting outcomes (e.g., favoring cost-effective winning projects vs. popular ones vs. winning projects with proportional representation of citizens’ preferences). These traits are extracted from feedback surveys of voters, who also cast ballots in the voting scenarios and therefore the survey data are linked to the actual voting behavior (see field study design in Section 4.2).
@@
    Voting alternatives: The three real-world voting scenarios determine the voting alternatives given as input to the AI representatives. They are illustrated in detail in Section 4.2. The actual City Idea scenario comes with 33 project alternatives and their cost, while the mock scenario comes with 5 costed project alternatives. The US elections comes with two alternatives. These three scenarios provide a broad spectrum of decisions spaces to assess the AI representatives.
@@
    Ballot formats: The two participatory budgeting scenarios are either approvals (n-approval ballot where ‘n’ is the number of projects approved) or cardinal preferences (score or cumulative) [48, 49]. In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest. In the case of cumulative ballots, the voters distribute a number of given points among the projects. Single-choice ballots are derived from single approvals, top scores or top points distributed to a project.
Generative AI. We use proprietary models such as GPT 3.5, GPT3 [50] and the open-source Llama2-13B [51]. GPT 3.5 and GPT 4 have similar performance for these tasks [52], hence we resort to GPT 3.5 for the analysis. Furthermore, both GPT 3.5 and GPT 4 use the “Let’s think step by step” to evaluate a prompt and find a solution that mimics human reasoning [52]. For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1), however the predictions with all models have the highest consistency to human choices, which we use for the analysis.
@@
Predictive AI benchmark. The personal human traits are set as features to predict the ballots using classification methods [53]. We sample 40% of the data and use synthetic minority oversampling [54] to augment the data. This 40% is decided empirically based on the performance of the neural network models. We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot. This model is then used to predict 60% of the unknown ballots and 40% of the known ballots.

### v4

@@ 2.2 AI representatives to recover from low voters turnout
AI representation of abstaining voters is more effective than arbitrary voters under fair collective choice. Figure 4 illustrates the capability of AI representatives to recover the consistency of voting outcomes lost by low voters turnouts. The three abstaining models (low engagement, trust, digital literacy) are assessed along with the baseline that determines random abstaining voters across the whole population. We emulate all four participation modalities as shown in Figure 1a. Strikingly, up to 75% of AI representation of low-engaged abstaining voters (94 representatives out of 126 abstaining voters in a population of 252 voters, see Section LABEL:sec:abstainingStats) is sufficient to recover up to 50% higher lost consistency than the random control population using equal shares. This superior consistency recovery is also observed for abstaining voters with low digital literacy (39.96%) and trust (25.44%). The fair aggregation rules of equal shares and Phragmén’s achieve on average 7.53% higher recovery compared to greedy for all the abstaining models. Even when controlling for the same number of winners, fair ballot aggregation methods achieve higher recovery than greedy by 6.65%. On average, AI representation by GPT-3.5 achieves higher recovery across all abstaining models, outperforming Llama3-8B by 4.3% and GPT-3 by 6.1% (Figures LABEL:fig:consistency-recovery_remainLABEL:fig:consistency-recovery-survey-USLABEL:fig:consistency-recovery_llamaLABEL:fig:consistency-recovery_GPT3, and Table LABEL:tab:false_positives_negatives_all_models).
Figure 4: Representing more than half of human abstaining voters with AI results in significant consistency recovery, in particular for fair ballot aggregation methods. Strikingly, AI representation of abstained voters is more effective than representing arbitrary voters (random control). Consistency recovery is at two levels: (i) False negative projects removed under abstaining but added back by AI representatives, which are are higher in ranking and number than (ii) false positive projects added under abstaining but removed by AI representatives. The consistency loss in voting outcomes by low voters turnout (x-axis) is emulated by removing different ratios of human voters (25%, 50%, 75% and 100%) among the whole population (baseline) and those who are likely to abstain: low engagement, trust and digital literacy profile (of the abstaining populations in the brackets on top). A consistency recovery (y-axis) is hypothesized by AI representation using GPT 3.5. (A) Actual participatory budgeting campaign of City Idea. (B) Studied participation modalities. (C)-(D) The origin of consistency recovery in participatory budgeting for utilitarian greedy and equal shares respectively. Abstaining voters result in falsely removing (left) and falsely adding (right) winning projects. AI representatives add back and remove these projects respectively to recover consistency. The projects and their probability to recover consistency under random control are shown for comparison.
AI representation of arbitrary vs. abstaining voters: from removing noise to restoring representation deficit. Figure 4c and 4d show the origin of inconsistency under utilitarian greedy and equal shares when voters abstain and how AI representatives recover from this. The figures show which projects are involved in consistency recovery and their ranking: (i) erroneously removed projects (false negatives, left) that are correctly added back by AI representatives (true positive) and (ii) erroneously added projects (false positive, right) that are correctly removed by AI representatives (true negative). Compared to true negatives projects, true positive ones are higher in ranking by an average of 7.2 and 2.5 positions for equal shares and utilitarian greedy respectively. The higher consistency recovery by the abstaining models compared to the random control population originates from an average of 0.71 and 0.47 additional projects involved in consistency recovery for the two ballot aggregation methods respectively. Moreover, the origin of consistency recovery by abstaining models is more prominent to true positive projects (mean of 1.66 over 1.0 for true negatives), while it is more prominent to true negative projects in the random control populations (mean of 2.27 over 1.89 for true positives). See Table LABEL:tab:recovery for a complete outline based on all the AI models. This result demonstrates a distinguishing quality of targeting the AI representation to abstaining voters: representation deficit is restored by adding back winners who would not be there otherwise, while a non-targeted AI representation has a noise-removal effect by removing erroneous winners. The district wise consistency recovery for the Aarau has been enumerated in Table LABEL:table:region.
2.3 Biases explaining AI (in)consistencies in choice and preference transitivity
Unraveling biases that explain AI inconsistencies. Figure 5 illustrates the biases that explain the (in)consistency of human-AI choice and the AI transitivity among different ballot formats (single choice vs. cumulative). We mainly show the results of the actual participatory budgeting of City Idea, while the results of the other datasets are shown in Figure LABEL:fig:ML_supple. We distinguish between (i) the inconsistencies originated by the three identified subpopulation of abstained voters and (ii) the inconsistencies by the AI representation of the entire population. Prediction models are constructed using recurrent neural networks (see Sections LABEL:sec:bias and LABEL:sec:ml), demonstrating robust performance with F1 scores averaging over 80% for abstaining groups and 74% for the entire population across all large language models (Table LABEL:tab:result_all). The relative importance of the personal human traits (independent variables) that explain the AI consistency for individual voters (dependent variable) is calculated using model agnostic Shapley additive explanations and local interpretable model-agnostic explanations [43] (see Section 4.4, Figures LABEL:fig:Lime_approvalLABEL:fig:Lime_score). For the abstaining subpopulations, the dependent variable is the consistency difference (plotted in Figure LABEL:fig:consistency-gain) between the one of abstaining voters and the mean of 10 random control subpopulations as shown in Section 2.2 to better isolate the biases within these subpopulations. For the AI representation of the entire population, the dependent variable is the choice consistency. To provide more robust evidence, we distinguish in Figures 5c and 5d those personal human traits that explain AI consistency (i) in all datasets, (ii) AI models and (iii) those which are statistically significant (p<0.05).
@@ 2.3 Biases explaining AI (in)consistencies in choice and preference transitivity
Figure 5: Compared to an arbitrary abstaining voter, those with low engagement and digital literacy exhibit characteristics that explain the consistency of human-AI representation and ballot formats, for instance, no interest in politics and support to education/health projects related to unconscious and surrogation biases. Time discounting, affect and conformity biases, such as preference for public space and environmental projects as well as support to families contribute to the consistency of human-AI choice. Time discounting factors such as preference for sport, and welfare projects as well as affect heuristics such as preference for projects that benefit families explain AI consistency among ballot formats. The relative importance of the personal human traits (y-axis) for the actual participatory budgeting campaign of City Idea over different large language models (x-axis) is depicted by the size of the bubbles and it is calculated using Shapley additive explanations. The consistency of human-AI representation (A) and ballot formats (B) (single choice vs. cumulative) is assessed. For each of these, the personal human traits explain the following: (i) The consistency difference between the three abstaining models and their random control. (ii) The (in)consistency of AI representation and transitivity for the whole population. The ‘-’ sign indicates non-significant values (p>0.05). (C)-(D) The statistically significant biases present in all AI models and datasets are summarized by chord diagrams.
Voters likely to abstain explain the consistency of their AI representation and transitivity of ballot formats. No interest in politics, support to self-benefit and children explain the consistency in AI representation of abstaining voters with low engagement compared to a random abstaining voter (Figure 5c): mean importance of 16.13% (p=0.04), 16.81% (p=0.04) and 9.32% (p=0.03) respectively. These traits relate to potential unconscious, egotism and affect biases [44, 45]. Potential surrogation and optimism biases stemming from support to education and public transit projects (mean importance of 17.23%, p=0.04 and 14.56%, p=0.04) explain the AI transitivity over different ballot formats for these abstaining voters (Figure 5d). The traits that explain the consistency in AI representation of abstaining voters with low digital literacy are support to city benefit (15.45%, p=0.02), children (8.11%, p=0.03), public transit (4.53%, p=0.04) and education (5.01%, p=0.02) that relate to altruism, conformity, optimism and surrogation biases (Figure 5c). Support to health (surrogation, 17.89%, p=0.03) and elderly (affect, 13.44%, p=0.03) projects explain the AI transitivity over different ballot formats for abstaining voters with low digital literacy (Figure 5d). These traits relate to older voters, who usually come with low digital literacy.
Affect and unconscious biases explain the (in)consistency of human-AI choice, while affect and time discounting biases explain transitivity of AI choice over ballot formats. The consistency of human-AI choice is explained by support to families (affect, 7.43%, p=0.03), public space (time discounting, 8.91%, p=0.01) and environment (conformity, 8.46%, p=0.03), while inconsistency is explained by support to elderly (affect, 11.39%, p=0.04). For the US elections, a political profile of left explains consistency of human-AI choice, while white voters explain inconsistency (33% higher than consistency, p<0.019, see Figure LABEL:fig:ML_supple). Affect and time discounting biases also explain the transitivity of AI representation over different ballot formats, in particular the support to families (18.22%, p=0.01), welfare (16.78%, p=0.02) and sport projects (17.05%, p=0.02). Figures 5LABEL:fig:Lime_scoreLABEL:fig:approvalLABEL:fig:Lime_approval, Table LABEL:tab:approval_tab and Section LABEL:sec:ml illustrate additional insights about how personal human traits explain the AI top choice and the human-AI consistency of the individual choices.
@@ 3 Discussion
The inevitable of generative AI voting and the race to safeguard democracy. Generative AI voting is likely to emerge as an inevitable technological convergence of AI and e-voting solutions currently adopted in real-world. On the one hand, generative AI and large language models are expected to become more open, pervasive and accessible to citizens [46, 15]. AI personal assistants are already part of everyday life [47, 48, 49], with their generative version expected to follow. On the other hand, the mandate of more direct, secure and active participation in decision making for public matters is expected to further scale up electronic voting solutions and digital platforms. For instance, participatory budgeting elections are mainly conducted digitally, while Estonia has already institutionalized a digital identity for 99% of its citizens as well as electronic voting since 2005 [50]. As the former president of Estonia emphasized "with the digital signature and the machine-readable ID card, we created the e-citizen". In the light of these converging technological advancements, the inter-operation of a generative personal voting assistant with a digital voting platform becomes technologically feasible along with the citizens’ need to have a more direct saying in several public matters and consultations. Therefore, the findings of this study become spot on to understand the implications of such a future, while they are significant to prepare timely safeguards for digital democracy.
@@
Fair voting design as a democratic safeguard to generative AI voting. Large language models prove not ready yet to accurately represent humans in complex voting scenarios such as participatory budgeting. We disentangle a complex interplay of localized and diverse context of the project alternatives, the large number of those and the underlying multi-faceted biases to which large language models are susceptible. However, a striking finding of this study, built upon our demonstrated real-world evidence in Aarau, is that consistency of large language models becomes significant under fair voting aggregation methods that promote a proportional representation of voters’ preferences. This motivates a huge opportunity to get democracy "right" in the digital era of AI: move to alternative voting methods that yield fairer voting outcomes for all, while shielding democratic outcomes from AI biases and inconsistencies. How to scale up these democratic blueprints remains though an open question. In particular, voting turnouts in participatory budgeting remain very low and far lower than other elections such as referenda or national elections. Despite eminent ethical and legal challenges of engaging AI representatives to democratic processes, our findings show that a more ethically-aligned AI representation of abstaining voters recovers consistency of voting outcomes, which would be lost in any case. This consistency loss can be to such large extent that is currently posing a long-standing barrier for participatory initiatives to take-off. Note that we focus on the AI representation of abstaining voters who intend to participate but their low engagement [41, 51], digital literacy [12, 52, 8, 53, 54] or trust [55, 56, 42] are a barrier for them. This is to distinguish voters whose abstaining is a conscious deliberate act and their AI representation would not be relevant or even desired in this context. Last but not least, our findings also demonstrate that AI representation alone does not suffice - a fair voting design is imperative to materialize significant recoveries from low voting turnouts.
@@
Why fair collective choice is resilient to AI biases and inconsistencies. We provide an explanation of this significant and striking finding. There is evidence that the method of equal shares has an inherent stability in the resulting voting outcomes [28]. Low and middle-cost projects require a very minimal support to get elected, and as a result, these winning projects are likely to be retained in the winning set, even with different choices or set of voters. Such projects are expected to be an origin of consistency. Indeed, this effect is also observed in the real-world voting scenario of City Idea, as with 80% abstaining voters, 84% of the winners are retained with equal shares, see Figure LABEL:fig:mixed_populations-abs and Table LABEL:table:change_gr_ES that shows the origin of this stability in terms of new projects added and removed in the winning set. Nevertheless, equal shares is still affected by low voters turnout, especially, for participation rates $<50\%$ [28]. As voters turnout in participatory budgeting is typically very low, these inconsistencies are both relevant and prevalent. Note that any comparison of stability between equal shares and utilitarian greedy should be made with caution as the number of winning projects under equal shares is much larger than utilitarian greedy. When we control for the number of winning projects between the two methods, equal shares remains more robust than utilitarian greedy, but to a much lower extent (see baseline [random] in Figure 4a and LABEL:fig:consistency-recovery_remain.)
Real-world testing of equal shares: overcoming a validity barrier. As City Idea promoted equal shares already in the project ideation phase and made apriori use of equal shares for the aggregation of the ballots, this study becomes the first of its kind with compelling realism and merit for the validity of our significant findings. This comes in stark contrast to other earlier studies [28, 27] that hypothesize the application of equal shares over proposed projects and ballots aggregated with the standard method of utilitarian greedy.
@@
Towards a trustworthy generative AI voting: a call for research and policy action. What information large language models use to reason about voting decisions is influential for different types biases to manifest. This is particularly the case for affect, unconscious and time discounting biases involved in AI representation of human choices and the transitivity of AI choices over different ballot formats. Abstaining voters with a low engagement, digital literacy and trust also possess related personal human traits that explain the consistency of their AI representation. Voters who come with a more active participation profile, without typical features of abstaining voters, appear irreplaceable as the reasoning of large language models cannot accurately estimate their choices (Figure LABEL:fig:consistency-recovery_opp). This motivates a tailored, purposeful and finite use of AI representation with the aim to set itself obsolete by recovering participation of abstaining voters, while mitigating for the consistency loss as long as voters abstain. Training data in generative AI voting are expected to play a key role for representative voting outcomes of the voters population. Ethical and democratic guidelines are urgently needed, particularly for the use of (generative) AI in voting processes. For instance, who shall determine the input training data of AI representatives? Should the training data involve only self-determined personal information of voters, or shall these be augmented with more universal knowledge and experts’ opinions? How to protect the privacy and autonomy of voters when training such AI representatives [49]? Will citizens retain power to control AI representatives that reflect their values and beliefs, while remaining accountable? These are some key questions as a basis of a call for action on research and policy in an emerging era of generative AI voting.
@@ 4.1 Emulating AI representatives
We introduce a generalized design to generate choices (ballots) and emulate AI voting representation using three large language models and a predictive AI method as a benchmark (referred to as ML).
Input for ballot generation. This comprises of (i) defining the voting scenario, the voting alternatives and the personal human traits of the voter as the input context (LLMs) or features (ML) to inform the ballot choice, (ii) generating the ballots and (iii) aggregating the ballots into a collective decision (see Figure 1).
@@
    Personal human traits: Voter’s information is used to define an AI persona to represent human voter, i.e. the AI representative (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2 in SI). This includes (i) socio-demographic characteristics (e.g., gender, age, education, household size), (ii) political interests (e.g., ideological profile, political belief), (iii) personal attitude towards project preferences (e.g., prioritization for green initiatives, sustainable transport, elderly care facilities) and (iv) preferences for the qualities of voting outcomes (e.g., favoring cost-effective winning projects vs. popular ones vs. winning projects with proportional representation of citizens’ preferences). These traits are extracted from feedback surveys of voters, who also cast ballots in the voting scenarios and therefore the survey data are linked to the actual voting behavior (Tables LABEL:table:prologue1LABEL:table:prologue2LABEL:table:prologue3LABEL:table:prologue4LABEL:table:prologue5 in Section 4.2).
@@
    Voting alternatives: The three real-world voting scenarios determine the voting alternatives given as input to the AI representatives. They are illustrated in detail in Section 4.2. The actual City Idea scenario includes 33 project alternatives and their cost, while the survey scenario comes with 5 costed project alternatives. The US elections however is associated with two alternatives. These three scenarios provide a broad spectrum of decisions spaces to assess the AI representatives.
@@
    Ballot formats: The two participatory budgeting scenarios are either approvals (n-approval ballot where ‘n’ is the number of projects approved) or cardinal preferences (score or cumulative) [57, 58]. In the case of score ballots, the voters provide a score to a project from a specified range such as 1 to 5 with 1 being the lowest preference and 5 being the highest. In the case of cumulative ballots, the voters distribute a number of given points (i.e. 10) among the projects. Single-choice ballots are derived from single approvals, top scores or top points distributed to a project.
Generative AI. We use proprietary models such as GPT 3.5, GPT 3 [59] and the open-source Llama3-8B [60]. For each of the large language models, predictions are made based on the five combinations of personal human traits (see Figure 1b), however the predictions with all models have the highest consistency to human choices, which we use for the analysis. We explored advanced OpenAI models, including GPT-4 Turbo, for the survey scenario of the City Idea campaign across various input context combinations related to personal human traits. The responses obtained under identical hyperparameter settings closely resemble those of GPT-3.5, with mean score deviations of only $\pm$0.05 across all alternatives. Since both GPT-4 and GPT-3.5 employ the same prompt evaluation logic for reasoning tasks [61] for standard prompt engineering techniques, we opted for the more cost-effective GPT-3.5 model for all prompt context variations and the different datasets.
@@
Predictive AI benchmark. The personal human traits are set as features to predict the ballots using classification methods [62]. Neural networks are used with 3 hidden layers and the relu [63] activation function for the sub-networks in the neural network. We employ the ensemble method [64] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [65]. This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data (more details in Section LABEL:sec:ml). We sample 40% of the data and use synthetic minority oversampling [66] to create synthetic data for augmenting the dataset. We conclude that 40% is the optimal proportion based on a standard approach of conducting multiple experiments with varying ratios of synthetic and real data [67, 68] and analyzing their impact on model performance. We also divide the prediction into multiple sub-problems, which predict the approvals or assigned scores for every option and then compile the final ballot. This model is then used to predict 60% of the unknown ballots and 40% of the known ballots.